# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ 4.1 Modeling Inter-Variable Dependencies
Spatial dependencies, or inter-time series relationships, play a pivotal role in affecting a model’s forecasting capability [28]. When presented with time series data and corresponding graph structures that delineate the strength of interconnections between time series, current studies typically employ (1) spectral GNNs, (2) spatial GNNs, or (3) a hybrid of both to model these spatial dependencies. At a high level, these methods all draw upon the principles of graph signal processing (as detailed in Definition 5 and subsequent discussion). Considering input variables $\mathbf{X}_{t}$ and $\mathbf{A}_{t}$ at a given time $t$, the goal here is to devise an effective GNN-based model, termed $\textsc{Spatial}(\cdot)$, to adeptly capture salient patterns between data points from different time series at time $t$. This can be expressed as $\hat{\mathbf{X}}_{t}=\textsc{Spatial}(\mathbf{X}_{t},\mathbf{A}_{t})$, where $\hat{\mathbf{X}}_{t}$ collects all time series representations at time $t$ with spatial dependencies embedded.
@@
Spectral GNN-Based Approaches. Early GNN-based forecasting models predominantly utilized ChebConv [108] to approximate graph convolution with Chebyshev polynomials, thereby modeling inter-time series dependencies. For instance, STGCN [66] intersects temporal convolution [109] and ChebConv layers to capture both spatial and temporal patterns. StemGNN [62] further proposes spectral-temporal graph neural networks to extract rich time series patterns by leveraging ChebConv and frequency-domain convolution neural networks. Other relevant research has largely followed suit, employing ChebConv to model spatial time series dependencies, while introducing innovative modifications. These include attention mechanisms [84, 92], multi-graph construction [54, 101], and combinations of the two [96]. Recently, building upon StemGNN, Jin et al. [28] have theoretically demonstrated the benefits of using spectral GNNs to model different signed time series relations, such as strongly positive and negative correlated variables within a multivariate time series. They also observed that any orthonormal family of polynomials could achieve comparable expressive power for such tasks, albeit with varying convergence rates and empirical performances.
@@
Spatial GNN-Based Approaches. Inspired by the recent success of spatial GNNs [29], another line of research has been modeling inter-time series dependencies using message passing [110] or graph diffusion [111, 112]. From the graph perspective, these methods are certain simplifications compared to those based on spectral GNNs, where strong local homophilies are emphasised [28, 113]. Early methods such as DCRNN [80] and Graph WaveNet [85] incorporated graph diffusion layers into GRU [114] or temporal convolution to model time series data. Subsequent works including GTS [64] and ST-GDN [115] also applied graph diffusion. In contrast, STGCN$(1^{st})$ (a second version of STGCN [66]) and ST-MetaNet [82] modelled spatial dependencies with GCN [116] and GAT [117] to aggregate information from adjacent time series. Related works, such as MRA-BGCN [86], STGNN*[87], GMAN[88], and AGCRN [91], proposed variants to model inter-time series relations based on message passing. To enhance learning capabilities, STSGCN [90] proposed spatial-temporal synchronous graph convolution, extending GCN to model spatial and temporal dependencies on localized spatial-temporal graphs. STFGNN [57] constructed spatial-temporal fusion graphs based on dynamic time wrapping (DTW) before applying graph and temporal convolutions. Z-GCNETs [95] enhanced existing methods with salient time-conditioned topological information, specifically zigzag persistence images. METRO [103] introduced multi-scale temporal graphs to characterize dynamic spatial and temporal interactions in time series data, together with the single-scale graph update and cross-scale graph fusion modules to unify the modeling of spatial-temporal dependencies. Another line of improvements incorporates graph propagation, allowing for the mixing of neighborhood information from different hops to learn high-order relations and substructures in the graph. Example are MTGNN [61] and SGP [104]. In particular, SGP exploits reservoir computing and multi-hop spatial processing to precompute spatio-temporal representations yielding effective, yet scalable predictive models. Follow-up works such as MTGODE [22] and TPGNN [97] proposed continuous graph propagation and graph propagation based on temporal polynomial coefficients. Other similar works include STGODE [79] and STG-NCDE [98]. Distinct from GAT-based methods, approaches based on graph transformer [118] can capture long-range spatial dependencies due to their global receptive field, making them a separate branch of enhanced methods. Examples include STAR [93], ASTTN [119], and ASTTGN [120]. At last, we mention NGAR [83] and GEN [94] as works that can predict the graph topology alongside node-level signals.
@@
Hybrid Approaches. Some hybrid methodologies also exist, integrating both spectral and spatial GNNs. For instance, SLCNN [89] employs ChebConv [108] and localized message passing as global and local convolutions to capture spatial relations at multiple granularities. Conversely, Auto-STGNN [121] integrates neural architecture search to identify high-performance GNN-based forecasting models. In this approach, various GNN instantiations, such as ChebConv, GCN [116], and STSGCN [90], can be simultaneously implemented in different spatial-temporal blocks.

### v3

@@ 4.1 Modeling Inter-variable Dependencies
Spatial dependencies, or inter-variable relationships, play a pivotal role in affecting a model’s forecasting capability [28]. When presented with time series data and corresponding graph structures that delineate the strength of interconnections between time series, current studies typically employ (1) spectral GNNs, (2) spatial GNNs, or (3) a hybrid of both to model these spatial dependencies. At a high level, these methods all draw upon the principles of graph signal processing (as detailed in Definition 5 and subsequent discussion). Considering input variables $\mathbf{X}_{t}$ and $\mathbf{A}_{t}$ at a given time $t$, the goal here is to devise an effective GNN-based model $\textsc{Spatial}(\cdot)$ to adeptly capture salient patterns from different time series. This can be expressed as $\hat{\mathbf{X}}_{t}=\textsc{Spatial}(\mathbf{X}_{t},\mathbf{A}_{t})$, where $\hat{\mathbf{X}}_{t}$ collects all time series representations at time $t$ with spatial dependencies embedded.
@@
Spectral GNN-based Approaches. Early GNN-based forecasting models predominantly utilized ChebConv [113] to approximate graph convolution with Chebyshev polynomials, thereby modeling inter-variable dependencies. For instance, STGCN [54] intersects temporal convolution [114] and ChebConv layers to capture both spatial and temporal patterns. StemGNN [50] further proposes spectral-temporal graph neural networks to extract rich time series patterns by leveraging ChebConv and frequency-domain convolution neural networks. Other relevant research has largely followed suit, employing ChebConv to model spatial time series dependencies, while introducing novel modifications. These include attention mechanisms [86, 105], multi-graph construction [41, 72], and combinations of the two [88].
@@
Spatial GNN-based Approaches. Inspired by the recent success of spatial GNNs [29], another line of research has been modeling inter-variable dependencies using message passing [115] or graph diffusion [116]. From the graph perspective, these methods are certain simplifications compared to those based on spectral GNNs, where strong local homophilies are emphasised [28, 117]. Early methods such as DCRNN [68] and Graph WaveNet [48] incorporated graph diffusion layers into GRU [118] or temporal convolution to model time series data. In contrast, STGCN$(1^{st})$ (a second version of STGCN [54]) and ST-MetaNet [76] modeled spatial dependencies with GCN [119] and GAT [120] to aggregate information from adjacent time series. To enhance learning capabilities, several improvements have been proposed. For example, STSGCN [92] proposed spatial-temporal synchronous graph convolution, extending GCN to model spatial and temporal dependencies on localized spatial-temporal graphs. MTGODE [23] and TPGNN [60] proposed continuous graph propagation and graph propagation based on temporal polynomial coefficients. Additionally, recent approaches based on graph transformer [100] or hypergraphs [121], such as [122] and [110], can capture longer-range spatial dependencies due to their global receptive field, making them a separate branch of enhanced methods.
@@
Hybrid Approaches. Some hybrid models also exist, integrating both spectral and spatial GNNs. For instance, SLCNN [95] employs ChebConv and localized message passing as global and local convolutions to capture spatial relations at multiple granularities. Conversely, Auto-STGNN [123] integrates neural architecture search to identify high-performance GNN-based forecasting models.
More details about modeling inter-variable dependencies in GNNs for time series forecasting are in Appendix A.1.